# The Impact of Baryonic Effects on the Dynamical Masses inferred using Satellite Kinematics

### v1

@@ A.2 Variations in profile of satellites and dark matter concentration
The middle panel of the bottom row of Fig. A.1 compares the ratios $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$ for four different combinations of $\mathcal{R}$, $\gamma$, and $\beta$, as indicated. This includes, in addition to our fiducial model with $\mathcal{R}=2$, $\gamma=1$, and $\beta=0$, a model with $\gamma=\mathcal{R}=1$ for which satellite galaxies trace the density distribution of the dark matter, a model with $\gamma=0$, for which $n_{\rm sat}(r)$ is cored, and a model with radially anisotropic kinematics for the satellites with $\beta=0.5$. Together, these choices span the range of satellite distributions found in observational studies (e.g. Carlberg et al. 1997; van der Marel et al. 2000; More et al. 2009a; Cacciato et al. 2013; Guo et al. 2012). Finally, the bottom-right panel shows the impact of changing the concentration-mass relation of the host halos. In particular, we increase or decrease the concentration of each halo by an amount that is representative of the scatter in the concentration-mass relation at fixed mass ($\sim 0.16$ dex), that is, we set $\log c_{200}\to\log c_{200}\pm 0.16$. Although each of these changes can strongly affect the shape of the velocity dispersion profile itself, they have negligible impact on the ratio of aperture velocity dispersions between the baryonic and DMO models.
Figure A.1: The aperture velocity dispersion ratio, $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$, as a function of halo mass, while varying many parameters in the halo model (see text). In each panel, the fiducial model is shown in black. If no baryonic corrections were applied to $\sigma_{\rm ap}$, the ratio would remain at $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}=1$, indicated with the black dotted line in each panel. The gray area indicates the extreme scenarios discussed in Section 4.3. Notably, all variations fall within this region, showing the minor impact of these parameters compared to these extreme scenarios.
@@ A.3 Redshift
Finally, we consider variations in the redshift of the primary $z$. Throughout the main text, we have adopted $z=0.1$, which is representative of the mean redshift of the SDSS spectroscopic sample. The bottom-left panel of Fig. A.1 compares these fiducial results to those obtained for $z=0.034$ and $z=0.184$, which are the extremes of the redshift range considered by Mitra et al. (2024).
@@
The redshift has a subtle impact on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. At lower $z$, the fibre collision scale of $5\arcseccorresponds$ to smaller physical scales, causing the increase of $\sigma_{\rm los}$ in the central region to start competing with the decrease in the outskirts. This is effectively the same as allowing $R_{\rm min}$ to extent further inwards (see Section 4.4). This effect is most pronounced in halos in which the stellar mass of galaxies is relatively dominant ($M_{\rm DMO}\sim 10^{12.5}\,\>{\rm M_{\odot}}$), where it results in an approximate 0.5% increase in $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$ for $z=0.034$. The curve for $z=0.184$ does not reach these halo mass scales because, at that redshift, $R_{\rm min}$ exceeds $R_{\rm max}$. Yet again, all these changes are minor in that they remain well within the gray-shaded region.

### v2

@@ A.2 Variations in profile of satellites and dark matter concentration
The left panel of the bottom row of Fig. A.1 compares the ratios $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$ for four different combinations of $\mathcal{R}$, $\gamma$, and $\beta$, as indicated. This includes, in addition to our fiducial model with $\mathcal{R}=2$, $\gamma=1$, and $\beta=0$, a model with $\gamma=\mathcal{R}=1$ for which satellite galaxies trace the density distribution of the dark matter, a model with $\gamma=0$, for which $n_{\rm sat}(r)$ is cored, and a model with radially anisotropic kinematics for the satellites with $\beta=0.5$. Together, these choices span the range of satellite distributions found in observational studies (e.g. Carlberg et al. 1997; van der Marel et al. 2000; More et al. 2009a; Cacciato et al. 2013; Guo et al. 2012). The middle-right panel in the bottom row shows the impact of changing the concentration-mass relation of the host halos. In particular, we increase or decrease the concentration of each halo by an amount that is representative of the scatter in the concentration-mass relation at fixed mass ($\sim 0.16$ dex), that is, we set $\log c_{200}\to\log c_{200}\pm 0.16$. Although each of these changes can strongly affect the shape of the velocity dispersion profile itself, they have negligible impact on the ratio of aperture velocity dispersions between the baryonic and DMO models. This principle is generic and therefore should also hold for more realistic radial profiles of the velocity anisotropy that are in better agreement with either observations (Mamon et al. 2019) and/or numerical simulations (Ascasibar and Gottlöber 2008; Mamon et al. 2010; Lemze et al. 2012; Lotz et al. 2019; van den Bosch et al. 2019).
Figure A.1: The aperture velocity dispersion ratio, $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$, as a function of halo mass, while varying many parameters in the halo model (see text). In each panel, the fiducial model is shown in black. If no baryonic corrections were applied to $\sigma_{\rm ap}$, the ratio would remain at $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}=1$, indicated with the black dotted line in each panel. The gray area indicates the extreme scenarios discussed in Section 4.3 and Fig. 5. Notably, all variations fall within this region, showing the minor impact of these parameters compared to these extreme scenarios.
@@ A.3 Redshift
Finally, we consider variations in the redshift of the primary $z$. Throughout the main text, we have adopted $z=0.1$, which is representative of the mean redshift of the SDSS spectroscopic sample. The bottom-right panel of Fig. A.1 compares these fiducial results to those obtained for $z=0.034$ and $z=0.184$, which are the extremes of the redshift range considered by Mitra et al. (2024).
@@
The redshift has a subtle impact on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. At lower $z$, the fibre collision scale of $5\arcseccorresponds$ to smaller physical scales, causing the increase of $\sigma_{\rm los}$ in the central region to start competing with the decrease in the outskirts. This is effectively the same as allowing $R_{\rm min}$ to extent further inward (see Section 4.4). This effect is most pronounced in halos in which the stellar mass of galaxies is relatively dominant ($M_{\rm DMO}\sim 10^{12.5}\,\>{\rm M_{\odot}}$), where it results in an approximate 0.5% increase in $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$ for $z=0.034$. The curve for $z=0.184$ does not reach these halo mass scales because, at that redshift, $R_{\rm min}$ exceeds $R_{\rm max}$. Yet again, all these changes are minor in that they remain well within the gray-shaded region.
Appendix B The effect of baryons on the kurtosis profiles
The primary observable considered in this work is the line-of-sight velocity dispersion profile, i.e., the second velocity moment. We additionally consider the fourth velocity moment of the line-of-sight velocity distribution, as implemented in Basilisk   (Mitra et al. 2024), which provides complementary dynamical information and helps break the mass–anisotropy degeneracy.
The projected fourth moment of the line-of-sight velocity dispersion is given by (e.g. Łokas 2002)
$\overline{v_{\rm los}^{4}}(R_{\rm p})=$ $\frac{2}{\Sigma(R_{\rm p})}\int_{R_{\rm p}}^{r_{\rm sp}}\left[1-2\beta\frac{R_{\rm p}^{2}}{r^{2}}+\frac{1}{2}\beta(1+\beta)\frac{R_{\rm p}^{4}}{r^{4}}\right]\,$
$\times\overline{v_{\rm r}^{4}}(r)n_{\rm sat}(r)\frac{r\mathrm{d}r}{\sqrt{r^{2}-R_{\rm p}^{2}}}\,.$ (B.1)
with
$\overline{v_{\rm r}^{4}}(r)=\frac{3G}{r^{2\beta}n_{\rm sat}(r)}\int_{r}^{r_{\rm s}}r^{2\beta}n_{\rm sat}(r)\sigma_{\rm r}^{2}(r)\frac{M(r)}{r^{2}}\mathrm{d}r.$ (B.2)
The corresponding line-of-sight kurtosis profile is then defined as
$\kappa_{\rm los}(R_{\rm p})=\frac{\overline{v_{\rm los}^{4}}(R_{\rm p})}{\sigma_{\rm los}^{4}(R_{\rm p})}.$ (B.3)
In these expressions, the radial velocity dispersion $\sigma_{\rm r}(r)$ and enclosed mass profile $M(r)$ are those described and varied throughout the main text.
Figure B.1 illustrates the results for the fiducial model. From left to right, we show the line-of-sight velocity dispersion profiles $\sigma_{\rm los}(R_{\rm p})$, identical to those shown in Fig. 2, the projected fourth moment of the line-of-sight velocity dispersion $\overline{v_{\rm los}^{4}}(R_{\rm p})$, and the corresponding line-of-sight kurtosis profiles $\kappa_{\rm los}(R_{\rm p})$. Results are shown for three halo masses, $M_{200}=10^{12.5}\,\>{\rm M_{\odot}}$ (top), $10^{13.5}\,\>{\rm M_{\odot}}$ (middle), and $M_{200}=10^{14.5}\,\>{\rm M_{\odot}}$ (bottom). We adopt the same baseline parameters as in the main text, $z=0.1$, $\mathcal{R}=2$, $\gamma=1$, and $\beta=0$, and we show no halo response ($\nu=0$) and adiabatic response ($\nu=1$). The shaded grey regions indicate the projected radial range $R_{\rm min}\leq R_{\rm p}\leq R_{\rm max}$ used in the Basilisk  analysis.
To interpret the behavior of the fiducial model, it is useful to consider the impact of the individual baryonic components in isolation (not shown). We first consider baryonic ejection by itself. In the extreme ejection limit, all baryons are removed and only 84% of the mass remains in dark matter. In this case, the total halo mass is reduced but the shape of the gravitational potential is unchanged. While this lowers the normalization of the line-of-sight velocity dispersion significantly, it leaves the ratio of moments, and thus the kurtosis profile, unchanged compared to the DMO case.
Next, we consider the gas component. If baryons are retained entirely as gas ($\sim$16% of the mass) with a cored profile transitioning to NFW, both $\sigma_{\rm los}$ and $\overline{v_{\rm los}^{4}}$ are slightly reduced relative to DMO. However, the suppression of the fourth moment is marginally weaker than the suppression of the variance squared, leading to a very slight increase in kurtosis. This effect, however, is very small. Likewise, any changes in the parameters of the diffuse component are negligible.
In contrast, the inclusion of a central galaxy significantly alters the gravitational potential in the inner regions ($R_{\rm p}\lesssim 0.1r_{200}$). The concentrated mass of the central galaxy boosts both $\sigma_{\rm los}$ and $\overline{v_{\rm los}^{4}}$ but the increase in the denominator ($\sigma_{\mathrm{los}}^{4}$) is much stronger than in the numerator. This leads to a strong suppression of the line-of-sight kurtosis at small radii. This effect dominates over all other baryonic components and drives the effects seen in the right side of the panels in Figure B.1. Its impact is most pronounced for lower-mass halos ($M_{\rm DMO}\sim 10^{12.5}\>{\rm M_{\odot}}$), where the stellar component is relatively more important, and becomes weaker at higher halo masses.
The adiabatic halo response ($\nu=1$, green curves) further amplifies this effect. The dark matter halo contracts and further deepens the central potential. This enhances the velocity dispersion in the center, further reducing the kurtosis in the inner regions.
Crucially, however, at larger projected radii where the influence of the central galaxy decreases, the kurtosis profile of the fiducial model converges to that of the DMO case. Consequently, across all models explored, differences in the kurtosis profile are negligible over the projected radial range $R_{\mathrm{min}}\leq R_{\mathrm{p}}\leq R_{\mathrm{max}}$ relevant for our satellite kinematics analysis.
Figure B.1: The impact of baryons on the fourth-order moment of satellite kinematics. We show results for three halo masses: $M_{\mathrm{DMO}}=10^{12.5}M_{\odot}$ (top), $10^{13.5}M_{\odot}$ (middle), and $10^{14.5}M_{\odot}$ (bottom). The panels display the line-of-sight velocity dispersion $\sigma_{\mathrm{los}}$ (left), the projected fourth velocity moment $\langle v_{\mathrm{los}}^{4}\rangle$ (middle), and the line-of-sight kurtosis $\kappa_{\mathrm{los}}\equiv\langle v_{\mathrm{los}}^{4}\rangle/\sigma_{\mathrm{los}}^{4}$ (right). Black curves represent the DMO model. Blue and green curves show the fiducial baryonic model without ($\nu=0$) and with ($\nu=1$) adiabatic halo response, respectively. While the central galaxy strongly suppresses $\kappa_{\mathrm{los}}$ in the inner halo ($R_{\mathrm{p}}\lesssim 0.1r_{200}$), the baryonic models converge to the DMO kurtosis profile in the radial range probed by the satellite kinematics analysis ($R_{\rm min}<R_{\rm p}<R_{\rm max}$, gray shaded region).